OpenStudy (anonymous):

Please help fast!! I will give medal. Find the area of the triangle having the indicated sides and angle with B=72 degrees 30', a=105 and c=64?

OpenStudy (anonymous):

\(Area=\frac{1}{2}105\times 64\times \sin(72.30')\)

OpenStudy (anonymous):

are you really working with minutes ? then you need to convert 30 minutes to .5 and compute \[\frac{1}{2}\times 105\times 64\times \sin(72.5)\]

OpenStudy (anonymous):

Is it 3200.94???

OpenStudy (anonymous):

i think i got it...is it 3204.49

OpenStudy (anonymous):

don't use 72.3 use 72.5 because calculator works with decimals, not minutes and seconds (old school)

OpenStudy (anonymous):

thank you so much, :)

OpenStudy (anonymous):

yw